# DietGlance: Dietary Monitoring and Personalized Analysis at a Glance with Knowledge-Empowered AI Assistant

### v2

@@ 7.1.3 Enhance the Integration of Multi-Devices and Proactive Intervention
While our system relied on smart glasses and laptops (or mobile phones), as suggested by participants, integrating additional devices such as wrist-worn activity trackers could provide a more comprehensive understanding of users’ health. Physical activity, sleep patterns, and metabolic rates are critical factors in dietary analysis and recommendations. Incorporating multimodal signals from multiple devices would enable more accurate, personalized feedback. For example, combining dietary data with physical activity metrics could refine calorie and nutrient recommendations, enhancing the system’s overall utility.
@@
Furthermore, leveraging multi-devices for real-time feedback and proactive dietary intervention can significantly enhance user engagement and adherence to dietary goals. By enabling users to quickly correct or label information on their smartwatches, the system can be further enhanced, as discussed in Sec. 7.1.1. Smart glasses or watches can deliver in-situ visual or audio cues [13] during meals, helping users make informed choices in a more seamless way. Real-time audio haptic feedback [75], for example, could combine sound cues (such as a pleasant tone indicating a correct portion or a warning beep for overeating) with tactile sensations (like vibrations to alert users when a meal exceeds recommended nutrient levels). This approach may not only improve user experience but also provide more intuitive and proactive interventions for dietary self-regulation based on the user’s current activity levels and dietary patterns, fostering healthier behaviors.
7.2 Generalizability
Our system could be applied to various scenarios in health and beyond by applying a similar analysis framework. While developed with Aria Glasses, our system can integrate with other eyewear for ingestive episode detection and supports non-wearable alternatives, allowing users to upload meal photos for analysis when wearables are unavailable or impractical. It can also be customized for regional diets by incorporating localized nutritional data, making it adaptable to diverse culinary preferences. Beyond dietary monitoring, the system can integrate fitness tracker data for physical activity analysis, assist with student tasks in education, and facilitate patient rehabilitation, where automated tracking and domain-specific analysis are essential.
7.3 Limitations and Future Work
@@
Eyewear Comfort
The Aria Glasses, though powerful, are heavy due to unnecessary sensors for our study. Participants reported discomfort from sweat, especially in summer. Many also required vision correction, but lens replacement was not practical. This discomfort may hinder long-term retention. Future research could explore lighter glasses, attachable sensors for personal eyewear, or alternatives like smart earbuds or watches.
@@
DietGlance accurately identified most foods but struggled with visually similar items. While DietGlance accurately identified most foods, it struggled with visually similar items. Users may need to manually update diet summaries. The system also assumed equal portions in shared meals, potentially leading to discrepancies. Additionally, estimating micronutrients like phosphorus and potassium was challenging due to variability in biological, environmental, and processing factors. Future work could focus on improving micronutrient datasets for higher accuracy. As emphasized by the experts, higher precise nutrient intake was expected in clinical contexts.
@@
The four-week duration of the longitudinal study is comparable with previous diet-related studies [9, 53, 71, 81, 66]. However, extending the study duration and incorporating additional metrics, such as cognitive load, would provide deeper insights into sustained engagement.
@@
Despite the system’s privacy-preserving consideration, participants still raised concerns, particularly in shared meal contexts. Clear communication protocols and transparent privacy practices for users and their companions can build trust and address these sensitivities. Future systems could also incorporate advanced safeguards against misinformation, such as prompt injection attacks [41] or erroneous inputs.
8 Conclusion

### v3

@@ 7.1.3. Enhance the Integration of Multi-Devices and Proactive Intervention
While our system relied on smart glasses and laptops (or mobile phones), as suggested by participants, integrating additional devices such as wrist-worn activity trackers could provide a more comprehensive understanding of users’ health (Jiang et al., 2023d, c). Physical activity, sleep patterns, and metabolic rates are critical factors in dietary analysis and recommendations. Incorporating multimodal signals from multiple devices would enable more accurate, personalized feedback (Jiang et al., 2023b). For example, combining dietary data with physical activity metrics could refine calorie and nutrient recommendations, enhancing the system’s overall utility.
@@
Furthermore, leveraging multiple devices for real-time feedback and proactive dietary intervention can significantly enhance user engagement and adherence to dietary goals. By enabling users to quickly correct or label information on their smartwatches, the system can be further enhanced, as discussed in Sec. 7.1.1. Smart glasses or watches can deliver in-situ visual or audio cues (Bressa et al., 2022) during meals, helping users make informed choices in a more seamless way. Real-time audio haptic feedback (Tan et al., 2024), for example, could combine sound cues (such as a pleasant tone indicating a correct portion or a warning beep for overeating) with tactile sensations (like vibrations to alert users when a meal exceeds recommended nutrient levels). This approach may not only improve user experience but also provide more intuitive and proactive interventions for dietary self-regulation based on the user’s current activity levels and dietary patterns, fostering healthier behaviors.
7.2. Generalizability
Our system’s analysis framework is designed to be generalizable across different hardware, regions, and domains.
While developed with Aria Glasses, our system can integrate with other eyewear equipped with similar sensors (e.g., IMU or cameras) for ingestive episode detection and supports non-wearable alternatives, allowing users to upload meal photos for analysis when wearables are unavailable or impractical.
While developed with Aria Glasses, our system supports integration with other eyewear equipped with similar sensors (e.g., IMU or cameras) for ingestive episode detection. Furthermore, to enhance accessibility, it supports non-wearable alternatives, allowing users to upload meal photos from any camera-enabled device for analysis when wearables are unavailable or impractical.
As our studies were conducted in Hong Kong, applying the current model directly to other regional diets might yield variable performance. However, the system’s overall pipeline is designed for modular adaptability. To extend DietGlance to other culinary contexts, users can update the knowledge base by replacing the nutritional database with a more region-specific source, narrow the retrieval scope by ingesting relevant local dietary guidelines into the RAG vector store, and adjust the system prompt to make it more specific (such as adding ‘cuisine type’). This modularity ensures the system can be swiftly swapped to cater to diverse global dietary standards.
Beyond dietary monitoring, the system’s framework has the potential to be extended to other health and wellness domains. With further development, such as developing APIs for data ingestion or updating the knowledge base, it could integrate fitness tracker data for physical activity analysis, assist with student tasks in education, or facilitate patient rehabilitation, where automated tracking and domain-specific analysis are essential.
7.3. Limitations and Future Work
@@
Eyewear Comfort and Practicality
A primary limitation is the practical burden and comfort of the eyewear. The Aria Glasses, though powerful, are heavy due to unnecessary sensors for our study. Participants reported discomfort from sweat, especially in summer. Many also required vision correction, but lens replacement was not practical. This discomfort may hinder long-term retention. Beyond physical comfort, the social acceptability of wearing smart glasses at every meal, especially in public or with companions, remains a realistic barrier to adoption. However, we observe the increasing commercialization and popularity of AI glasses by Meta, Rokid, etc., which suggests that these devices are becoming more normalized and that social acceptability may improve as they become more commonplace. Future research could explore lighter glasses, attachable sensors for personal eyewear, or alternatives like smart earbuds or watches.
@@
DietGlance accurately identified most foods but struggled with visually similar items. Users may need to manually update diet summaries. The system also assumed equal portions in shared meals, potentially leading to discrepancies. Additionally, estimating micronutrients like phosphorus and potassium was challenging due to variability in biological, environmental, and processing factors. This aligns with recent findings that noninvasive micronutrient assessment remains a major open challenge in the field (Balch et al., 2025). Future work could focus on improving micronutrient datasets for higher accuracy. As emphasized by the experts, more precise nutrient intake was expected in clinical contexts. Specifically, this could involve developing an agentic framework that proactively prompts the user for clarification on ‘hidden’ factors (e.g., ‘Did you add salt or soy sauce?’) when high-uncertainty foods are detected. Additionally, integrating GPS-based context retrieval (with user consent) could allow the system to fetch precise nutritional data from restaurant-specific menus rather than generic databases.
Limitation of RAG Module
As discussed in Sec.  7.1.2, while our RAG module significantly improved factual accuracy, its high dependency on the quality of the retrieved documents also introduced new challenges, such as the curation of the knowledge base. Moreover, other methods like fine-tuning could be explored to improve the performance, potentially by learning a more practical and actionable style. Therefore, another promising direction for future work is the exploration of agentic models to intelligently decide when to retrieve verifiable facts via RAG and when to rely on fine-tuned internal knowledge to better balance accuracy with actionability.
@@
The four-week duration of the longitudinal study is comparable with previous diet-related studies (Biel et al., 2018; Morshed et al., 2022; Silva et al., 2023; Weinreich et al., 2017; Santas et al., 2012). However, extending the study duration and incorporating additional metrics, such as cognitive load, would provide deeper insights into sustained engagement.
@@
Despite the system’s privacy-preserving consideration, participants still raised concerns, particularly in shared meal contexts. Clear communication protocols and transparent privacy practices for users and their companions can build trust and address these sensitivities. Future systems could also incorporate advanced safeguards against misinformation, such as prompt injection attacks (Liu et al., 2023) or erroneous inputs.
Participant Demographics and Generalizability.
While our system showed promise in handling complex cuisines (e.g., Chinese-style food), our study population was primarily Chinese participants. Although they consumed both Western and Eastern foods, and our system can be adapted by incorporating localized nutritional data (as discussed in Sec. 7.2), this demographic homogeneity limits the generalizability of our findings on the behavioral patterns, cultural preferences, and social acceptability of the system. Future work is required to validate DietGlance with more diverse, multicultural participant pools.
Nutrition Library Limitations.
While we constructed our nutrition library from high-trust, expert-vetted sources, and the experiment results show that our RAG module significantly improved the accuracy of the system’s outputs compared to the baseline LLM, we did not perform an exhaustive, line-by-line check for minor inconsistencies between the documents. Future work could explore automated methods for consistency checking across the corpus, such as by using contradiction detection models to identify conflicting recommendations. Moreover, as new nutritional research is published, future systems could build a robust pipeline for dynamically updating and re-vetting the knowledge base to ensure the RAG module’s accuracy is maintained over time.
Database Dependency.
Our expert evaluation relied on a specific local nutritional database (Centre for Food Safety, Hong Kong) as the ground truth. We acknowledge that nutritional values can vary across different national databases due to regional food differences. Furthermore, our system integrates multiple databases to maximize food coverage; however, this introduces data heterogeneity that may affect perceived accuracy when evaluation is confined to a single regional standard. While we chose the most geographically relevant database for our expert evaluation, the system’s performance might vary if evaluated against different national standards, highlighting the importance of localizing knowledge bases for deployment. Future work may develop adaptive RAG frameworks that can automatically select and retrieve from the most regionally appropriate nutritional database based on the user’s location or cultural context.
8. Conclusion